OpenStudy (anonymous):

What is the slope of the line passing through the points (0, –5) and (4, 2)?

OpenStudy (anonymous):

\[slope= \frac{ y _{2} -y _{1}}{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so im guessing that i have to do 0-(-5)=0+5=5

ganeshie8 (ganeshie8):

(0, –5) and (4, 2) x1 y1 x2 y2

ganeshie8 (ganeshie8):

plugin now

OpenStudy (anonymous):

but its gonna be 5 but eaither -5 or 5

OpenStudy (anonymous):

??

OpenStudy (anonymous):

so i got to do 4+2=6 and -5/6=-1.2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{ 2-(-5) }{ 4-0 }\]

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

so its 2-(-5)=2+5=7 and 4-0= 4 so its

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ty for your help :)

OpenStudy (anonymous):

u r welcome

OpenStudy (anonymous):

how many are you able to do? cuz i have one more

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but u should now how to solve it

OpenStudy (anonymous):

i mean learn not just take the answer

OpenStudy (anonymous):

What is the slope of the line whose equation is 2x – 4y = 10? -2 -1 2 4 and 1/2 or one half .